OpenStudy (darkbluechocobo):

An object with a mass of 5 kg has 9kJ of kinetic energy. Calculate the velocity.

OpenStudy (godlovesme):

\[k=1/2mv^2\] that's the formula ^^^^ can u try solving

OpenStudy (darkbluechocobo):

so V=1/2(5)9^2?

OpenStudy (godlovesme):

yes :)

OpenStudy (darkbluechocobo):

202.5?

OpenStudy (godlovesme):

 u can round it

OpenStudy (darkbluechocobo):

man I messed up S:

OpenStudy (godlovesme):

it's fine :) but do u see what i did?

OpenStudy (darkbluechocobo):

You did not move the v around a k is the kinetic energy so I just screwed up my equation then you divided by 2.5 t oget to both sides then sqrt it to get v^2 to just v

OpenStudy (godlovesme):

exactly =)

OpenStudy (darkbluechocobo):

hmm can you go over another practice problem to see if I can correct my mistake

OpenStudy (godlovesme):

sure!

OpenStudy (darkbluechocobo):

So this is one is: A 98N object travels at 10m/s. What is the kinetic energy.

OpenStudy (darkbluechocobo):

So K=1/2(98)(10^2)?

OpenStudy (darkbluechocobo):

so 1/2(98)=49

OpenStudy (godlovesme):

so far so good but what about the 10^2 ?

OpenStudy (darkbluechocobo):

100*49=4900

OpenStudy (godlovesme):

yep good job :)

OpenStudy (darkbluechocobo):

would that be the final answer?

OpenStudy (godlovesme):

yes

OpenStudy (darkbluechocobo):

S: There is no answer for 4900 there is an option for 490

OpenStudy (godlovesme):

:O is it 490 ki

OpenStudy (darkbluechocobo):

I would assume it

OpenStudy (godlovesme):

we forgot to change the newton to kg

OpenStudy (darkbluechocobo):

ohhh

OpenStudy (darkbluechocobo):

so would that explaiin it then?

OpenStudy (godlovesme):

may be :P